OpenStudy (anonymous):

solve the system of equations x+y+z=13 -2x-y+z=-4 x-2y-z=-18

OpenStudy (anonymous):

x+y+z=13 -2x-y+z=-4 Add. -3x + 2z = 9 Then. (-2x-y+z=-4)(times) -2 4x+2y-2z=8 x-2y-z=-18 Add. 5x - 3z = -10 Now (-3x + 2z = 9)(times)3 -9x +6z = 27 (5x - 3z = -10)(times)2 10x - 6z = -20 ~~~~~~~~~~~~~~~ -9x +6z = 27 10x - 6z = -20 Add. x = 7 ~~~~~~~~~~~~~~~~~ Now substitute x in 5x - 3z = -10 5(7) - 3z = -10 35 - 3z = -10 Subtract 35 on both sides -3z = -45 Divide by -3 on both sides z = 15 ~~~~~~~~~~~~~~~~~~~~~~ Now substitute x and z in the first equation. 7 + y + 15 = 13 Combine like terms 22 + y = 13 Subtract 22 on both sides y = -9 ~~~~~~~~~~~~~~~~~ Substitute all variables in every equation to see if they are correct. x = 7 y = -9 z = 15 7+(-9)+15=13 7-9=-2+15=13 13=13 Correct! -2(7)-(-9)+15=-4 -14+9=-5+15=10 10=-4 Incorrect!! 7-2(-9)-15=-18 7+18=25-15=10 10=-18 Incorrect!! Not sure where I went wrong.... @Best_Mathematician Can you help me find where I went wrong?

OpenStudy (anonymous):

@jim_thompson5910 Help me find my mistake?

jimthompson5910 (jim_thompson5910):

when you add equations 1 and 2 you should get x+y+z=13 -2x-y+z=-4 ------------ -1x+0y+2z=9 which turns into -x+2z = 9

OpenStudy (anonymous):

Ohhhh! Duh!

OpenStudy (anonymous):

looks like he/she found what she want! still need help...just tag me in comment

OpenStudy (anonymous):

x+y+z=13 -2x-y+z=-4 Add. -x + 2z = 9 Then. (-2x-y+z=-4)(times) -2 4x+2y-2z=8 x-2y-z=-18 Add. 5x - 3z = -10 Now (-x + 2z = 9)(times)3 -3x +6z = 27 (5x - 3z = -10)(times)2 10x - 6z = -20 ~~~~~~~~~~~~~~~ -3x +6z = 27 10x - 6z = -20 Add. 7x = 7 x=1 ~~~~~~~~~~~~~~~~~ Now substitute x in 5x - 3z = -10 5(1) - 3z = -10 5 - 3z = -10 Subtract 5 on both sides -3z = -15 Divide by -3 on both sides z = 5 ~~~~~~~~~~~~~~~~~~~~~~ Now substitute x and z in the first equation. 1 + y + 5 = 13 Combine like terms 6 + y = 13 Subtract 6 on both sides y = 7 ~~~~~~~~~~~~~~~~~ Substitute all variables in every equation to see if they are correct. x = 1 y = 7 z = 5 1+7+5=13 8+13=13 13=13 Correct! -2(1)-7+5=-4 -2-7=-9+5=-4 -4=-4 Correct! 1-2(7)-5=-18 1-14=-13-5=-18 -18=-18 Correct!